The Chicago Mercantile Exchange (CME) E-mini S&P 500 index futures contract for December 2021 (symbol: ESZ1) closed at 3845.00. Suppose that the S&P 500 index closed at 3875.00, and that the interest rate is 2%. What is the implied dividend yield on the index?


1
Expert's answer
2021-03-30T07:04:07-0400

"F=" "3845.00, S=" "3875.00, r=2\%"

F=S+INTEREST SAVED - DIVIDEND FOREGONE

"3845=3875+2\% of 3875-dividend"

"dividend=3875+2\%of 3875-3845=3875+77.5-3845=107.5"

dividend yield"=dividend/S=107.5/3875=0.027742=2.7742\%"